# From Slices to Structures: Unsupervised 3D Reconstruction of Female Pelvic Anatomy from Freehand Transvaginal Ultrasound

### v1

@@ 4.4 Parameter Initialization
  • On-slice initialization: When the spatial positions and orientations of the US slices are known (e.g., via a tracking system), we distribute the Gaussian centers uniformly across the slices. This ensures that Gaussians are positioned exactly where they are needed from the start, resulting in strong gradient signals and faster convergence.
  • Uniform sampling in a bounding volume: In cases where slice poses are unknown or unreliable, we sample Gaussian centers uniformly from a predefined 3D bounding box that encloses the volume of interest.
While this strategy is more agnostic and flexible, it often requires additional optimization time for Gaussians to migrate toward active regions. The full mathematical formulation and a visual comparison of both methods is provided in Sec. .2.
@@ 4.5 Inactive Gaussians and Density Control
Such inactive Gaussians waste computational resources during both the forward and backward passes. To mitigate this, we incorporate a lightweight yet effective density control mechanism that periodically removes such Gaussians and replaces them with new ones better positioned to contribute meaningfully.
@@
Gaussians for which $m_{g}<\epsilon$ (where $\epsilon$ is a small predefined threshold) are considered inactive. These are removed from the parameter set every $N$ optimization steps. This process is efficient to implement and avoids the overhead of evaluating gradient norms or rendering error sensitivity for Gaussians.
To maintain a constant number of Gaussians in the scene, we reintroduce new ones to replace those pruned. Rather than randomly sampling across the entire volume (which could repeat the problem), we restrict the new means $\mathbf{\mu}_{g}$ to be sampled uniformly within the current bounding box of active Gaussians:
$\mathbf{\mu}_{\text{new}}\sim\mathcal{U}[\mathbf{\mu}_{\min},\,\mathbf{\mu}_{\max}]$
This simple heuristic leverages the existing distribution of active Gaussians to guide new candidates toward potentially informative regions, without requiring a full analysis of image-space reconstruction error. Other parameters are reset to default values as described in Sec. 4.4.
As a complementary strategy, we apply a non-uniform learning rate schedule where the Gaussian means $\mathbf{\mu}_{g}$ are updated more aggressively than other parameters (e.g., opacity, intensity, and rotation). The motivation is to allow Gaussians to quickly migrate toward regions where they influence the rendered image before their visibility is ”shut off” by the optimizer. In more extreme variants, we temporarily freeze non-positional parameters entirely during early training epochs. This prevents premature deactivation and leads to more effective exploration of the volume.
Both the pruning–reinitialization mechanism and learning rate scheduling are essential for maintaining a healthy and responsive population of Gaussians during training. Full details are provided in Sec. .3.
@@ 4.6 Practical Considerations for Implementation
To make training tractable at scale, we apply a series of CUDA-level optimizations across both the forward and backward passes of the differentiable rasterizer. In particular, we precompute invariant quantities such as inverse covariance matrices ($\Sigma_{g}^{-1}$), which can then be reused across all affected pixels. We also adopt shared memory tiling, where Gaussian parameters are loaded once per tile and reused by all threads in a block, thereby minimizing global memory traffic. Further efficiency is gained through symmetry-aware matrix handling, which exploits the symmetric structure of covariance matrices to reduce both memory usage and compute overhead, and through matrix-free computation strategies that replace expensive general-purpose matrix operations with lightweight element-wise arithmetic. Finally, compiler-level tuning, including the use of aggressive optimization flags and fast math approximations, further accelerates execution. Together, these optimizations improve training throughput by an order of magnitude compared to a naive implementation and are critical for scaling to tens of thousands of Gaussians in high-resolution 3D volumes. Detailed descriptions and performance results are provided in Sec. .5.
@@ 5.1 Experimental Setup
We conducted initial experiments to identify optimal default parameters for training. Image resolution and batch size were selected to balance reconstruction quality with computational constraints, particularly GPU memory. The remaining settings were empirically tuned to ensure high-fidelity outputs across diverse inputs. Unless otherwise specified, all subsequent experiments use the configuration detailed in Table 2. The selection of specific values is explained in more detail in Sec. 5.2. We divide the experiments into optimizations (Sec. 5.2) and reconstruction quality (Sec. 5.3).
Table 2: Default configuration for all experiments.
Category Value
Input resolution $\times256$ px
Slices per sweep 100
Number of Gaussians 120 per slice ($\approx$ 12k)
Loss Hybrid L1 (80%) + SSIM (20%)
Learning rates Means:0.20, Opacity:0.03, Scale:0.01, Intensity:0.008
Batch size 32 slices
Epochs 60
To comprehensively evaluate model performance, we report standard image similarity metrics: Structural Similarity Index Measure (SSIM), Mean Absolute Error (MAE), Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Normalized Cross-Correlation (NCC), using their conventional formulations.
@@ 5.2 Optimizations
As outlined in Sec. 4, applying Gaussian Splatting to TVS introduces specific difficulties due to the shift from a pinhole camera model to the physical characteristics of US image formation. Key issues include identifying a suitable loss function, defining effective initialization strategies, and addressing common problems such as inactive Gaussians. In the following, we detail the solutions adopted to address these.
5.2.1 Training Objective
Selecting an appropriate loss function is essential, as it shapes the optimization trajectory and affects both the accuracy and perceptual quality of the final reconstruction. Each loss emphasizes different aspects of fidelity—for instance, local pixel differences or broader structural consistency. To determine the most effective objective for US image reconstruction within our framework, we assessed a range of commonly used losses: L1, L2, SSIM, PSNR, and NCC. We also tested hybrid configurations, referred to as “Hybrid L1” (80% L1, 20% SSIM) and “Hybrid SSIM” (20% L1, 80% SSIM). To ensure a fair comparison, we first determine the optimal learning rate for each objective through preliminary experiments without advanced optimizations. After determining the optimal rates, we trained the model under each loss configuration and assessed performance across a broad set of evaluation metrics. A quantitative comparison of the results is provided in Table 3. Some qualitative results are in Fig. 5. Showing robustness in most metrics, the hybrid L1 loss is selected for further optimizations and experiments.
Table 3: Metric values for evaluated loss functions. Hybrid L1 denotes 80% L1 and 20% SSIM, whereas Hybrid SSIM indicates 20% L1 and 80% SSIM.
Loss Type L1 ($\downarrow$) L2 ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) NCC ($\uparrow$)
Hybrid SSIM 0.0192 0.0016 0.8663 27.86 0.9791
Hybrid L1 0.0188 0.0016 0.8528 28.07 0.9799
L1 0.0198 0.0017 0.8197 27.72 0.9782
L2 0.0245 0.0023 0.7651 26.41 0.9699
NCC 3.4030 33.7651 0.4314 -15.28 0.9762
PSNR 0.0210 0.0016 0.8048 27.88 0.9792
SSIM 0.0198 0.0017 0.8695 27.61 0.9792
Figure 5: Images rendered using commonly used loss functions, such as L1, L2, SSIM, PSNR, and NCC. We also tested hybrid configurations, referred to as “Hybrid L1” (80% L1, 20% SSIM) and “Hybrid SSIM” (20% L1, 80% SSIM).
@@ 5.2.2 Results of Optimization Techniques
To evaluate the suitability of each of the initialization approaches described in Sec. 4.4, we run experiments and report the L1 and SSIM for each case, as shown in Table 4. Fig. 6 shows an example of the resulting scene obtained through uniform random initialization.
Table 4: L1 and SSIM metrics on runs with different initialization for sagittal-only, transversal-only and sagittal & transversal view optimization. Number of Gaussians: 1̃2k
Group Init. method L1 ($\downarrow$) SSIM ($\uparrow$)
Sagittal Random 0.0067 0.9610
On-Slice 0.0055 0.9694
diff. 0.0012 0.0084
Transversal Random 0.0047 0.9717
On-Slice 0.0037 0.9783
diff. 0.0010 0.0066
Both Random 0.0060 0.9655
On-Slice 0.0054 0.9697
diff. 0.0006 0.0042
Figure 6: Example of a 3D view scene after uniform random initialization of means.
Table 4 shows that on-slice initialization yields slightly better results, and is therefore chosen as the default mean initialization method
@@
The total number of Gaussian primitives used to represent the volume is a key hyperparameter that directly influences both reconstruction quality and computational cost. Increasing this number provides the model with more flexibility to capture fine anatomical details, potentially enhancing reconstruction fidelity. However, this improvement comes with a higher memory footprint and significantly longer training times.
To explore this trade-off, we systematically evaluated the method’s performance by varying the number of Gaussians on a fixed set of US slices (specifically, real data sweeps). We monitored reconstruction quality metric alongside the total training time, as shown in Fig. 7. As expected, a higher number of Gaussians generally leads to improved reconstruction quality, reflected in lower L1 loss and higher SSIM scores. At the same time, training time increases substantially due to the added computational burden per iteration.
Figure 7: L1 Loss and SSIM vs. training time for different numbers of Gaussians representing the scene.
For the tested 256x256 resolution images, the results suggest an optimal range between approximately 60,000 Gaussians (about 278 per slice in the test setup) and 100,000 Gaussians (about 463 per slice), which appears to strike a good balance between reconstruction quality and computational efficiency. Beyond roughly 100,000 Gaussians, the gains in L1 and SSIM begin to plateau, indicating diminishing returns, while the execution time continues to rise significantly. To strike a good balance between performance and execution time during experiments, we use approximately 12k Gaussians.
@@
As described before, inactive Gaussians can limit the effect of a large number of primitives on the final reconstruction. We now present the results for both parameter-specific learning rates and density control.
Parameter-specific Learning Rates: We perform an exhaustive parameter search, obtaining the results shown in Table 2. A more comprehensive analysis and specific results can be found in Sec. .3.
Density Control: We apply reinitialization of Gaussians with spatial awareness, as described in Sec. 4.5, using different thresholds for pruning ($m_{g}=|I_{g}o_{g}|$ values of 0.2, 0.1, 0.05, and 0.01) and tested several schedules for when to apply the removal and addition steps. The best results were obtained using a threshold of 0.05 and applying density control every 10 epochs, starting at epoch 10 and continuing until epoch 50 (specifically at epochs 10, 20, 30, and 40, given the 60-epoch training setup).
Fig. 8 shows how this method impacted training compared to the baseline without mitigation. In the L1 loss plot, the red line (“With” density control) shows spikes after each control step, caused by abrupt changes in the set of Gaussians. The loss then stabilizes, and by the end of training, both L1 and SSIM slightly outperform the baseline without density control.
Figure 8: Training performance with (red) and without (blue) density control, compared to the baseline using default learning rates. The spikes in the red L1 curve mark the moments when density control was applied.
Fig. 9 shows the intensity, opacity, and scale distributions for the baseline (no mitigation of inactive Gaussians) and for the two proposed techniques. Notably, intensity values can be negative, allowing the model to offset overly bright areas by subtracting signal during composition. Without mitigation, distributions collapse into consistent patterns: intensity clusters around zero, opacities shift lower, and scales often fall below the initial 1.65, reducing each primitive’s spatial impact. In contrast, both parameter-specific learning rates and density control yield broader distributions.
Figure 9: Parameter distributions for no mitigation of inactive Gaussians (left), parameter-specific learning rates (middle) and density control (right).
Finally, when combining the two presented techniques to handle inactive Gaussians, we did not observe significant improvement over using them separately. For this reason, we adopt parameter-specific learning rates as the default technique for Gaussian deactivation in our approach.
@@ 5.3.1 View-Specific Reconstructions on Synthetic Data
Based on insights gained from earlier experiments, we now evaluate the reconstruction performance of our method more closely. We begin with synthetic data, using the following setup, number of epochs set to 60, image resolution is $\times256$ pixels, number of Gaussians per slice is 276, and total number of slices is 100. With this configuration, training on sagittal and transversal sweeps took about 5 minutes each. The resulting L1 losses were 0.0049 for sagittal and 0.0034 for transversal views, with SSIM scores of 0.972 and 0.980, respectively. Fig. 10 shows the corresponding reconstructions. We observed cone-shaped artifacts when using models trained on one view to reconstruct the other. These are caused by the angle distribution of slices, which span from -60° to 60° around their respective axes. Visually, the reconstructions are quite faithful, although small inaccuracies appear near the edges, likely due to abrupt intensity transitions in the ground truth masks. Since the synthetic data represents segmentation masks, generating 3D reconstructions is straightforward. Fig. 11 displays the 3D models reconstructed from sagittal-only, transversal-only, and combined sweeps. Combining both views yields the most complete result, but even a single sweep captures a good amount of structural information.

### v2

@@ 4.3 Parameter Initialization
- On-slice initialization: Given a set of slice poses—whether inferred via our sensorless motion model or obtained from external tracking—we distribute Gaussian centers uniformly across the image planes. This places primitives directly within the signal domain from the start, ensuring strong initial gradient updates and accelerating convergence.
- Uniform sampling in a bounding volume: Alternatively, we sample Gaussian centers uniformly within a predefined 3D bounding box enclosing the volume of interest. While this strategy is agnostic to the initial slice configuration, it typically requires additional optimization iterations for primitives to migrate from empty space toward informative anatomical regions.
@@ 4.4 Inactive Gaussians and Density Control
Such inactive Gaussians waste computational resources. To mitigate this, we incorporate a lightweight density control mechanism that periodically removes such Gaussians and replaces them with new ones better positioned to contribute meaningfully.
@@
Gaussians for which $m_{g}<\epsilon$ (where $\epsilon$ is a low predefined threshold) are considered inactive. These are removed from the parameter set every $N$ optimization steps. This process is efficient to implement and avoids the overhead of evaluating gradient norms or rendering error sensitivity for Gaussians.
To maintain volumetric capacity, we re-seed an equivalent number of new Gaussians, sampling their means uniformly within the bounding box of currently active primitives ($\mathbf{\mu}_{\text{new}}\sim\mathcal{U}[\mathbf{\mu}_{\min},\,\mathbf{\mu}_{\max}]$). This simple heuristic leverages the existing distribution of active Gaussians to guide new candidates toward potentially informative regions, without requiring a full analysis of image-space reconstruction error. Other parameters are reset to default values as described in Sec. 4.3.
As a complementary strategy, we apply a non-uniform learning rate schedule where the Gaussian means $\mathbf{\mu}_{g}$ are updated more aggressively than other parameters (e.g., opacity, intensity, and rotation). The motivation is to allow Gaussians to quickly migrate toward regions where they influence the rendered image before their visibility can be “shut off” by the optimizer. In more extreme variants, we temporarily freeze non-positional parameters entirely during early training epochs. This prevents premature deactivation and leads to more effective exploration of the volume.
Both the pruning–reinitialization mechanism and learning rate scheduling are essential for maintaining a healthy and responsive population of Gaussians during training.
4.5 Joint Pose and Scene Optimization
A core challenge in sensorless freehand ultrasound is the variability in probe motion. While we initialize our reconstruction assuming a smooth, linear angular sweep, human hand motion inevitably introduces irregularities. To address this, we decouple the slice poses from the fixed geometric prior. We treat the pose parameters $\mathbf{y}_{i}=[\mathbf{r}_{i},\mathbf{t}_{i}]$ for each slice $i$ as learnable parameters. During the backward pass, gradients are propagated through the rasterizer not only to the Gaussian primitives but also to the transformation matrices defining the slice positions. This allows the model to fine-tune the alignment of individual slices to maximize photometric consistency with the reconstructed volume. We apply a lower learning rate to pose parameters compared to scene parameters to prevent drift and maintain the global structure defined by the initialization.
@@ 4.6 Practical Considerations for Implementation
To make training tractable at scale, we apply a series of CUDA-level optimizations across both the forward and backward passes of the differentiable rasterizer. In particular, we precompute invariant quantities such as inverse covariance matrices ($\Sigma_{g}^{-1}$), which can then be reused across all affected pixels. We also adopt shared memory tiling, where Gaussian parameters are loaded once per tile and reused by all threads in a block, thereby minimizing global memory traffic. Further efficiency is gained through symmetry-aware matrix handling, which exploits the symmetric structure of covariance matrices to reduce both memory usage and compute overhead, and through matrix-free computation strategies that replace expensive general-purpose matrix operations with lightweight element-wise arithmetic. Together, these optimizations improve training throughput by an order of magnitude compared to a naive implementation and are critical for scaling to hundreds of thousands of Gaussians in high-resolution 3D volumes.
@@ 5.1 Experimental Setup
We followed standard practices for image resolution and batch size, while empirically tuning remaining settings to ensure high fidelity. Unless otherwise specified, we train for 150 epochs with a batch size of 32. We use a resolution of $256\times 256$ for synthetic data and $256\times 128$ for real data, the latter selected to match the dataset’s native 2:1 form factor. The scenes are represented with approximately 50k Gaussians, and optimized using a hybrid loss ($\lambda_{\text{L1}}=0.8,\lambda_{\text{SSIM}}=0.2$) with learning rates 0.2 (means), 0.03 (opacity), 0.01 (scale), and 0.008 (intensity). The selection of specific values is explained in more detail in Sec. 5.2. We divide the experiments into optimizations (Sec. 5.2) and reconstruction quality (Sec. 5.3). To comprehensively evaluate model performance, we report standard image similarity metrics: Structural Similarity Index Measure (SSIM), Peak Signal-to-Noise Ratio (PSNR), and Learned Perceptual Image Patch Similarity (LPIPS), using their conventional formulations.
@@ 5.2 Optimizations
In this section, we evaluate the impact of various method configurations on reconstruction performance. First, we explored different training objectives, testing L1, L2, SSIM, PSNR, NCC, LPIPS and hybrid combinations. We found that a hybrid configuration ($\lambda_{\text{L1}}=0.8,\lambda_{\text{SSIM}}=0.2$) yielded the best results, offering superior robustness across metrics compared to the other tested functions. Consequently, we select this loss for all subsequent experiments.
@@ 5.2.1 Results of Optimization Techniques
To evaluate the suitability of each of the initialization approaches described in Sec. 4.3, we run experiments and report the SSIM, PSNR and LPIPS for each case, as shown in Table 2. Fig. 5 shows an example of the resulting scene obtained through uniform random initialization.
Table 2: SSIM, PSNR and LPIPS metrics on runs with different initialization for sagittal and transversal view optimization.
Group Init. method SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS($\downarrow$)
Sagittal Random 0.960 26.49 0.083
On-Slice 0.975 28.54 0.057
Transversal Random 0.974 28.25 0.062
On-Slice 0.983 30.38 0.047
Figure 5: Example of a reconstructed 3D view scene after uniform random initialization of means.
As shown in Table 2, on-slice initialization yields consistently superior metrics and is selected as the default method. We attribute this performance gap to the sparse nature of US slice data. Random initialization frequently places primitives in empty space far from the image planes; these primitives contribute negligible density to the rendering, resulting in vanishing gradients that prevent them from migrating toward anatomical structures. In contrast, on-slice initialization guarantees that primitives spatially overlap with valid signal from the first iteration. This ensures immediate gradient flow and prevents primitives from becoming effectively “inactive” at the start of training—a challenge we analyze further in the following section on density control.
@@
The number of Gaussian primitives directly influences both reconstruction quality and computational cost. Increasing this number provides the model with more flexibility to capture fine anatomical details, potentially enhancing reconstruction fidelity. However, this improvement comes with a higher memory footprint and significantly longer training times.
To explore this trade-off, we systematically varied the number of Gaussians on a fixed set of real data sweeps. We monitored reconstruction quality metric alongside the total training time, as shown in Fig. 6. As expected, a higher number of Gaussians generally leads to improved reconstruction quality, reflected in lower L1 loss and higher SSIM scores. At the same time, training time increases substantially due to the added computational burden per iteration.
Figure 6: Mean SSIM, PSNR and LPIPS vs. training time for different numbers of Gaussians representing the scene.
For the tested 256x128 resolution images, the results suggest an optimal range between approximately 60,000 Gaussians (about 600 per slice in the test setup) and 100,000 Gaussians (about 1000 per slice), which appears to strike a good balance between reconstruction quality and computational efficiency. Beyond roughly 100,000 Gaussians, the gains in L1 and SSIM begin to plateau, indicating diminishing returns, while the execution time continues to rise significantly. To strike a good balance between performance and execution time during experiments, we use approximately 50k Gaussians.
@@
As described before, inactive Gaussians can limit the effect of a large number of primitives on the final reconstruction. We apply reinitialization of Gaussians with spatial awareness, as described in Sec. 4.4, using different thresholds for pruning ($m_{g}=|I_{g}o_{g}|$ values of 0.2, 0.1, 0.05, and 0.01) and test several schedules for when to apply the removal and addition steps. The best results were obtained using a threshold of 0.05 and applying density control every 25 epochs, starting at epoch 50 and continuing until epoch 125.
We observed that this mechanism introduces transient spikes in the loss function immediately following each pruning and reseeding step, though the optimization stabilizes rapidly. Quantitatively, the strategy yields minor improvements, primarily when using random initialization where the risk of inactive primitives is higher. In this setting, applying density control improved SSIM from 0.974 to 0.978, PSNR from 28.25 dB to 28.82 dB, and LPIPS from 0.062 to 0.053. Thus, while not strictly necessary for convergence—especially given the robustness of our default on-slice initialization—it serves as a useful refinement to maximize performance.
5.2.2 Robustness to Motion Jitter
Real-world sweeps often contain “jitter”—small deviations in speed and angle due to hand tremor or irregular motion. To evaluate our method’s robustness, we introduced synthetic jitter to the simulated dataset. We define jitter as random noise added to the ground truth poses, scaled by a percentage of the total sweep extent (e.g., 2.5% jitter corresponds to deviations up to 2.5% of the total rotation range and image dimensions).
Since pose optimization is an optional component of our pipeline (Sec. 4.5), we study it as an ablation to assess how it improves robustness to motion jitter. Quantitative results for increasing jitter levels are summarized in Table 3.
As expected, introducing motion jitter degrades the performance of the fixed-pose baseline. At 2.5% jitter—a level consistent with a reasonably steady hand—the test SSIM drops from 0.971 to 0.931.
Activating learnable poses allows the model to correct for these deviations. Pose optimization significantly recovers performance, boosting Test SSIM from 0.931 back to 0.952, increasing PSNR and reducing LPIPS error. This demonstrates that our method can self-calibrate small motion artifacts without external tracking.
At extreme jitter levels (5%), the optimization struggles to recover the correct geometry, likely getting trapped in local minima where the misalignment is too large for the gradients to guide the slices back to coherence. Interestingly, at this high noise level, the fixed-pose baseline outperforms the optimized version on the test set. This suggests that for highly erratic motion, the strong regularization of the smooth geometric prior is safer than attempting to optimize poses. However, for standard clinical sweeps, joint optimization provides a significant boost in fidelity.
Table 3: Effect of learned poses on data affected by different percentages of jitter.
Training Test
SSIM PSNR LPIPS SSIM PSNR LPIPS
No Jitter
TVGS 0.978 28.36 0.053 0.972 28.51 0.058
TVGS w/o pose opt. 0.975 28.54 0.057 0.971 28.83 0.059
2.5% Jitter
TVGS 0.950 23.06 0.099 0.952 23.88 0.089
TVGS w/o pose opt. 0.942 21.13 0.106 0.931 20.44 0.126
5% Jitter
TVGS 0.927 19.47 0.122 0.899 18.62 0.159
TVGS w/o pose opt. 0.923 19.06 0.126 0.913 19.12 0.145
@@ 5.3.1 View-Specific Reconstructions on Synthetic Data
Based on insights gained from earlier experiments, we now evaluate the reconstruction performance of our method more closely. We first evaluate our method on the synthetic dataset to validate geometric accuracy in a controlled environment. Using the default configuration of approximately 50,000 Gaussians, the optimization converges rapidly, requiring approximately 8 minutes for each view (sagittal and transversal).
Qualitative Results. Fig. 7 illustrates the reconstruction fidelity. The model successfully captures the scene geometry, producing visually faithful renderings. We observe minor artifacts near the edges of structures, which we attribute to the sharp, non-smooth intensity transitions characteristic of binary segmentation masks used to generate the synthetic ground truth. Fig. 8 displays the resulting 3D volumetric representations. While individual sagittal or transversal sweeps capture the general structure, they exhibit cone-shaped artifacts due to the limited angular coverage of the ultrasound probe (spanning $\pm 60$°). Combining both views minimizes these occlusions, resulting in a more complete 3D reconstruction.
Comparison with Implicit Representations. Benchmarking unsupervised 3D US reconstruction is inherently challenging due to the lack of standard baselines that operate without external tracking. We compare our approach against ImplicitVol [22], a state-of-the-art implicit neural representation method. To ensure a fair comparison in our tracking-free setting, we adapted ImplicitVol to use the slice poses estimated by our framework rather than pre-computed tracked poses.
Quantitative results are summarized in Table 4. Our method achieves high structural similarity and peak signal-to-noise ratios, confirming the validity of the reconstruction. We observe that ImplicitVol achieves marginally higher quantitative scores (SSIM and PSNR) and lower perceptual error (LPIPS). This likely stems from the continuous nature of the MLPs in ImplicitVol, possessing an inductive bias towards smoothness well-suited for perfect, noise-free synthetic surfaces. In contrast, our discrete Gaussian representation may introduce minor high-frequency noise when approximating these perfectly smooth boundaries.
However, this slight difference in metric performance is outweighed by the computational efficiency of our approach. As shown in Table 4, ImplicitVol requires approximately one hour to converge. In contrast, our Gaussian Splatting framework achieves comparable fidelity in just 8 to 16 minutes—representing an improvement in training speed of nearly an order of magnitude.
@@
Table 4: Quantitative comparison of our method with different number of Gaussians against ImplicitVol on simulated data.
Train Test
SSIM PSNR LPIPS SSIM PSNR LPIPS Time
Sagittal
ImplicitVol 0.994 31.35 0.018 0.994 34.67 0.013 56m
TVGS (Ours) - 50k 0.978 28.36 0.053 0.972 28.51 0.058 8m
TVGS (Ours) - 100k 0.981 29.00 0.047 0.974 29.02 0.051 16m
Transversal
ImplicitVol 0.991 30.09 0.026 0.984 31.22 0.047 62m
TVGS (Ours) - 50k 0.983 30.33 0.046 0.970 28.25 0.065 8m
TVGS (Ours) - 100k 0.986 31.47 0.039 0.972 28.55 0.054 16m